OpenStudy (anonymous):

Find two consecutive integers whose sum is 73. Which of the following equations could be used to solve the problem? 2x = 73 2x + 1 = 73 2x + 2 = 73 x2 + 1= 73

OpenStudy (anonymous):

Tow consecutive integers can be represented by x and x+1. So write an equation.

OpenStudy (aggie12341):

2x+2=73

OpenStudy (anonymous):

2x+2=73 @Easyaspi314

OpenStudy (anonymous):

Why 2x+2? The sum of x and x+1 = x+ x+1 = 2x + 1.

OpenStudy (anonymous):

@Cweezie23 Why 2x + 2? What is the sum of x and x+1?

OpenStudy (anonymous):

@Cweezie23 Well?

OpenStudy (anonymous):

idk this is extra credit we havent learned yet so im tryin to raise my grade

OpenStudy (anonymous):

so whats the answer @Easyaspi314

OpenStudy (anonymous):

please

OpenStudy (anonymous):

x + (x+1) = 2x + 1 When you add x to x, you get 2x. so you have 2x + 1 = 73.

OpenStudy (anonymous):

Choose the correct mathematical translation for the given statement. The luggage can weigh no more than 70 pounds. x < 70 x ≤ 70 x > 70 x ≥ 70

OpenStudy (anonymous):

Two consecutive integers are integers one after another, like 23 and 24, or 567 and 568 So sisnce we do not know what these integers are, we cann the fitrst one x, and the second one which is 1 more than thr first one, we call it x+1. so we want the sum of the two integers to be 73. SUM = add the two numbers so, x + (x+1) = 73 2x + 1 = 73

OpenStudy (anonymous):

You're welcome.

OpenStudy (aggie12341):

2x+1=73